OpenStudy (anonymous):

points (2,2) and (5,5) lie on line l. what is the slope of the line that is parallel to l?

OpenStudy (luffingsails):

good morning.

OpenStudy (luffingsails):

So, if both points line on the line... find the slope between the two points. That will be the slope of the line but it will also be the slope of any line that is parallel to the line l too. Can you find the slope between the two points?

OpenStudy (anonymous):

I'm just on a Mastery test on my credit recovery at school currently and my class will be up soon, I'm not really familiar with this material and good morning ^^

OpenStudy (luffingsails):

To be able to calculate slope: slope = rise / run = (y1 - y2) / (x1 - x2) For this problem: slope = (5-2) / (5 - 2) = 3 / 3 = 1. That's your answer.

OpenStudy (anonymous):

thanks you very much ~

OpenStudy (luffingsails):

Welcome!

OpenStudy (anonymous):

would you mind helping with 4 more? that's all I need to complete this test

OpenStudy (anonymous):

if not I understand